OpenStudy (anonymous):

Of 6000 apples harvested, every third apple was too small, every fourth apple was too green, and every tenth apple was bruised. The remaining apples were perfect. How many apples were too small, too green, bruised, or perfect? What if n apples were harvested? @ganeshie8

ganeshie8 (ganeshie8):

`every third apple was too small,` since you have `6000` apples, number of small apples would be : \[\large \dfrac{6000}{3} \] yes ?

OpenStudy (anonymous):

I think so

ganeshie8 (ganeshie8):

simplifying gives you \(\large 2000\)

ganeshie8 (ganeshie8):

can you work # of green and bruised apples similarly?

OpenStudy (anonymous):

6000/4=1500 6000/10=600

ganeshie8 (ganeshie8):

yes!

ganeshie8 (ganeshie8):

what about perfect apples ?

ganeshie8 (ganeshie8):

and what about apples that are both too small and green ?

OpenStudy (anonymous):

I did 6000/3 =2000 , 6000/4= 1500 then 2000+1500= 3500

ganeshie8 (ganeshie8):

so you're adding all the defective apples ?

ganeshie8 (ganeshie8):

looks you forgot the bruised ones

OpenStudy (anonymous):

when adding them all it would be 4100

ganeshie8 (ganeshie8):

yes don't you want to subtract the apples that are both too `small and green` ?

ganeshie8 (ganeshie8):

how many apples are both too `small and green` ?

ganeshie8 (ganeshie8):

` every third apple was too small, every fourth apple was too green and every tenth apple was bruised. ` that means every 12th apple is both too small and green every 40th apple is both green and bruised every 30th apple is both brused and too small and every 60th apples all too small, green and bruised

ganeshie8 (ganeshie8):

Overall : number of defective apples without repitition would be \[\dfrac{6000}{3} + \dfrac{6000}{4} + \dfrac{6000}{10} - \dfrac{6000}{12} - \dfrac{6000}{40} - \dfrac{6000}{30} + \dfrac{6000}{60}\]

ganeshie8 (ganeshie8):

simplify and subtract that from 6000 for the number of good apples

OpenStudy (anonymous):

after doing the math that you put above I got 3150

ganeshie8 (ganeshie8):

yes! thats the count of total defective apples

ganeshie8 (ganeshie8):

subtract it from 6000 for the good apples

OpenStudy (anonymous):

After subtracting 6000-3150 I got 2850

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

I lady in my class said the answer is 2800 which the professor said could be right but I don't see how

ganeshie8 (ganeshie8):

your lady is right, try again : \[\dfrac{6000}{3} + \dfrac{6000}{4} + \dfrac{6000}{10} - \dfrac{6000}{12} - \dfrac{6000}{\color{Red}{20}} - \dfrac{6000}{30} + \dfrac{6000}{60}\]

OpenStudy (anonymous):

that doesn't add up either cause I got 3200

ganeshie8 (ganeshie8):

yes those are the defective apples

ganeshie8 (ganeshie8):

u still need to subtrct 3200 from 6000 for the good ones right?

OpenStudy (anonymous):

that works!

OpenStudy (anonymous):

how do I find out "What if n apples were harvested?"

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

For n apples harvested, rewrite all the fractions you have with n in the numerator instead of 6000.

OpenStudy (anonymous):

so instead of saying 6000/ by all of the numbers it would be n*(1/3+1/4+1/10-1/12-1/20-1/30+1/60)

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

okay thanks i understand what i have to do now i just wasn't sure